### v1

# A machine-learning photometric classifier for massive stars in nearby galaxies

Abstract
Context. Mass loss is a key aspect of stellar evolution, particularly in evolved massive stars, yet episodic mass loss remains poorly understood. To investigate this, we need evolved massive stellar populations across various galactic environments.
Aims. However, spectral classifications are challenging to obtain in large numbers, especially for distant galaxies. We addressed this by leveraging machine-learning techniques.
Methods. We combined Spitzer photometry and Pan-STARRS1 optical data to classify point sources in 26 galaxies within 5 Mpc, and a metallicity range 0.07-1.36 Z. Gaia DR3 astrometry was used to remove foreground sources. Classifications are derived using a machine-learning model developed by Maravelias et al. (2022).
Results. We report classifications for 1,147,650 sources, with 276,657 sources ($\sim 24\%$) being robust. Among these are 120,479 Red Supergiants (RSGs; $\sim 11\%$). The classifier performs well even at low metallicities ($\sim 0.1$ Z) and distances under 1.5 Mpc, with a slight decrease in accuracy beyond $\sim 3$ Mpc due to Spitzer ’s resolution limits. We also identified 21 luminous RSGs ($\textrm{log}(L/L_{\odot})\,$$\geq 5.5$), 159 dusty Yellow Hypergiants in M31 and M33, as well as 6 extreme RSGs ($\textrm{log}(L/L_{\odot})\,$$\geq 6$) in M31, challenging observed luminosity limits. Class trends with metallicity align with expectations, though biases exist.
Conclusions. This catalog serves as a valuable resource for individual-object studies and James Webb Space Telescope target selection. It enables follow-up on luminous RSGs and Yellow Hypergiants to refine our understanding of their evolutionary pathways. Additionally, we provide the largest spectroscopically confirmed catalog of massive stars and candidates to date, comprising 5,273 sources (including $\sim 330$ other objects).
Key Words.:
Stars: massive – Stars: mass-loss – Stars: evolution – Galaxies: individual: WLM, NGC 55, IC 10, M31, NGC 247, NGC 253, NGC 300, IC 1613, M33, Phoenix Dwarf, NGC 1313, NGC 2366, NGC 2403, M81, Sextans B, NGC 3109, NGC 3077, Sextans A, NGC 4214, NGC 4736, NGC 4826, M83, NGC 5253, NGC 6822, Pegasus DIG, NGC 7793 – Methods: statistical – Catalogs
1 Introduction
One of the main goals of James Webb Space Telescope (JWST) is to study galaxies over cosmic time, from the early Universe to now. Their light though is a combination of various components, one of which is their stellar content. And, although rare in absolute numbers, massive stars have an important contribution. Through their feedback, whether from strong stellar winds or explosive supernovae, massive stars play a critical role in enriching and shaping the environments of their host galaxies. This is especially important in the early Universe, when metallicity was extremely low and the formation and evolution of such stars is still not well understood. The only way to gain insight into these objects is by examining these populations in nearby low metallicity galaxies. Resolved population studies in such galaxies are possible in the Local Group (e.g. Sextans A; Lorenzo et al. 2022, LMC; Vink et al. 2023, SMC; Shenar et al. 2024) but challenging at larger distances. Therefore, we lack well-explored populations of massive stars at these metalliclities.
The main goal of the ASSESS [^1] http://assess.astro.noa.gr/ (”Episodic Mass Loss in Evolved Massive stars: Key to Understanding the Explosive Early Universe”) project (Bonanos et al., 2024) was to investigate the role of episodic mass loss in the evolution of massive stars (see e.g. Yang et al. 2023; Antoniadis et al. 2024, 2025; de Wit et al. 2024; Munoz-Sanchez et al. 2024a, b; Zapartas et al. 2024). A large number of sources with secure classifications was needed to explore its importance across various metallicity environments, therefore we set up both an observing campaign to acquire spectra for a large number of sources (see e.g. de Wit et al., 2023; Bonanos et al., 2024; de Wit et al., subm.) as well as a machine-learning approach (Maravelias et al. 2022; henceforth Paper I). We developed a classifier that uses optical and IR photometry to select dusty, mass losing, evolved massive stars and classify Spitzer detected point sources into different broad classes. The purpose of this classifier was to predict the classes for approximately 1.2 M sources from 26 galaxies within 5 Mpc and spanning a metallicity range (0.07-1.36 Z), creating the largest point-source catalog with spectral-type classifications for and beyond the Local Group. In Paper I, we provide the method and explore the classifier’s prediction accuracy [^2] Code and other material availabe at https://github.com/gmaravel/pc4mas., while the current paper presents the results of the classifier’s application and the corresponding catalog of sources.
In Section 2 we present the data collection and processing, the selection process of the best candidates (based on the results from the machine-learning classifier), as well as the collection of all sources with known spectral classification from literature. In Section 3 we describe the catalog we compiled along with statistics regarding the number of objects per class, and present color-magnitude diagrams (CMDs). In Section 4 we provide a comparison of the spectral types we predicted with the ones derived from the literature and discuss the performance of the classifier. We present the populations as a function of metallicity and explore luminous Red Supergiants (RSGs) and dusty Yellow Supergiants (YSGs) found in our sample. Finally, in Section 5 we summarize and conclude our work.
2 Data collection and processing
In the following sections, we describe our sample and the steps we followed to build our catalogs as well as to remove foreground sources. Although this has been extensively discussed in Paper I, we provide here a short overview, along with adjustments performed mainly regarding the foreground cleaning due to the release of Gaia DR3.
2.1 Galaxy sample
In Paper I we used M31 and M33 for training the classifier, and WLM, Sextans A, and IC 1613 for testing it. In this work, we present the results of applying the classifier to the whole sample of galaxies included in the ASSESS project (Bonanos et al., 2024) (except for the LMC and SMC which were treated separately, e.g. Yang et al. 2019, 2020, 2021, 2023; de Wit et al. 2023; Antoniadis et al. 2024; Munoz-Sanchez et al. 2024a). This is presented in Table 1 along with the basic properties of the galaxies, including their names, coordinates, types, radii (which correspond to the sizes of the galaxy based on visual inspection that was used to match the catalogs), distances, and metallicity.
Table 1: Galaxies examined in this work along with some basic properties.
Galaxy R.A. (J2000) Dec. (J2000) Typea Radius Distanceb Metallicityc
(hh:mm:ss) (dd:mm:ss) (′) (Mpc) ($Z_{\sun}$)
WLM 00:01:58 --15:27:39 SB(s)m: sp 9 0.95$\pm$0.01 0.131
NGC 55 00:14:54 --39:11:48 SB(s)m? edge-on 21 1.98$\pm$0.02 0.312
IC 10 00:20:17 +59:18:14 dIrr IV/BCD 6 0.78$\pm$0.04 0.473
M31 00:42:44 +41:16:09 SA(s)b LINER 105 0.75$\pm$0.02 1.364
NGC 247 00:47:09 --20:45:37 SAB(s)d 14 3.56$\pm$0.03 0.545
NGC 253 00:47:33 --25:17:18 SAB(s)c 21 3.61$\pm$0.03 0.836
NGC 300 00:54:53 --37:41:04 SA(s)d 15 1.94$\pm$0.04 0.417
IC 1613 01:04:48 +02:07:04 IB(s)m 14 0.72$\pm$0.01 0.168
M33 01:33:51 +30:39:37 SA(s)cd HII 30 0.85$\pm$0.02 0.659
Phoenix Dwarf 01:51:06 --44:26:41 IAm 8 0.43$\pm$0.01 0.0710
NGC 1313 03:18:16 --66:29:54 SB(s)d 8 4.21$\pm$0.06 0.3511
NGC 2366 07:28:55 +69:12:57 IB(s)m 6 3.21$\pm$0.04 0.1612
NGC 2403 07:36:51 +65:36:09 SAB(s)cd 14 3.13$\pm$0.06 0.5613
M81 09:55:33 +69:03:55 SA(s)ab 18 3.61$\pm$0.22 $0.60^{14}$
Sextans B 10:00:00 +05:19:56 IB(s)m 5 1.38$\pm$0.08 0.1015
NGC 3109 10:03:07 --26:09:35 SB(s)m edge-on 13 1.37$\pm$0.08 0.1216
NGC 3077 10:03:19 +68:44:02 I0 pec 5 3.75$\pm$0.11 0.8917
Sextans A 10:11:01 --04:41:34 IBm 4 1.41$\pm$0.05 0.0715
NGC 4214 12:15:39 +36:19:37 IAB(s)m 7 2.82$\pm$0.09 0.3218
NGC 4736 12:50:53 +41:07:14 (R)SA(r)ab 11 4.34$\pm$0.08 0.4819
NGC 4826 12:56:44 +21:40:59 (R)SA(rs)ab 7 4.36$\pm$0.06 0.2220
M83 13:37:01 --29:51:56 SAB(s)c 10 4.79$\pm$0.09 0.9721
NGC 5253 13:39:56 --31:38:24 Im pec 4 3.55$\pm$0.21 0.3722
NGC 6822 19:44:58 +14:48:12 IB(s)m 10 0.46$\pm$0.01 0.3023
Pegasus DIG 23:28:36 +14:44:35 dI 8 0.95$\pm$0.03 0.1024
NGC 7793 23:57:50 +32:35:28 SA(s)d 8 3.39$\pm$0.06 0.4825
[…]
1Urbaneja et al. (2008); 2Hartoog et al. (2012); 3Cosens et al. (2024); 4Zurita & Bresolin (2012, range 1.05-1.66 $Z_{\sun}$); 5Davidge 2021, similar to LMC, 0.43 Z from Hunter et al. (2007) and M33 - see below; 6Spinoglio et al. (2022); 7Kudritzki et al. (2008); 8Bresolin et al. (2007); 9U et al. (2009, range 0.3-1.0 $Z_{\sun}$); 10Ross et al. (2015); 11Hadfield & Crowther (2007); 12Thuan & Izotov (2005); 13Bresolin et al. (2022, range 0.35-0.77 $Z_{\sun}$); 14Arellano-Córdova et al. (2016, range 0.37-0.78 $Z_{\sun}$); 15Kniazev et al. (2005, range 0.07-0.14 $Z_{\sun}$); 16Hosek et al. (2014); 17Storchi-Bergmann et al. (1994); 18Pilyugin et al. (2015); 19Moustakas & Kennicutt (2006); 20Kang et al. (2020, from iron abundance, range 0.03-0.40 $Z_{\sun}$); 21Hernandez et al. (2019, range 0.35-1.58 $Z_{\sun}$); 22Monreal-Ibero et al. (2012); 23Patrick et al. (2015); 24Skillman et al. (1997, range 0.06-0.20 $Z_{\sun}$) 25Della Bruna et al. (2021).
333 \$a\$\$a\$footnotetext: Derived from NASA/IPAC Extragalactic Database (NED). \$b\$\$b\$footnotetext: From Tully et al. (2023), except for M81, Sextans B, NGC 3109, and NGC 5253 which were derived from Tully et al. (2013). \$b\$\$b\$footnotetext: Metallicities are based on young clusters or massive stars or H ii regions, when available. An average value was reported in cases where multiple measurements or radial gradients exist. For solar abundance we used 12+log(O/H) = 8.69 (Asplund et al., 2009).
Figure 1: Example of the fitting processing and foreground removal in NGC 2403. Top panel: Gaia sources where the green ellipse defines the boundary selected for the galaxy. Middle panel: Fitting the parallax data for all sources within NGC 2403. The foreground contribution (dark gray dotted line) is scaled according to the densities of sources inside and outside the galaxy, while the purple dashed line shows the (Gaussian) contribution of galactic sources, and the blue line indicates the total fit. The vertical black dashed lines show the $3\sigma$ limits defined by the Gaussian properties of the galactic sources. Bottom panel: Similar to parallax, but for proper motion in declination (see Section 2.3 for more details).
2.2 Surveys used and data processing
Since IR photometry has been proven successful in identifying evolved massive stars (Britavskiy et al., 2014, 2015, 2019; Kourniotis et al., 2017) and those with dusty environments (Bonanos et al., 2009, 2010, 2024), we based our catalogs on pre-compiled point-source catalogs from the Spitzer Space Telescope (Boyer et al. 2015; Khan et al. 2015; Khan 2017; Williams & Bonanos 2016 - see Table 2 for details; the Spitzer column provides the number of sources within the radius as presented in Table 1). The mid-IR photometry ($\rm 3.6\,\mu m$, $4.5\,\mu\rm m$, $5.8\,\mu\rm m$, $8.0\,\mu\rm m$, $24\,\mu\rm m$ [^4] We are using the [3.6], [4.5], [5.8], [8.0], and [24] notation in the remainder of the paper.) was cross-matched with optical photometry ($g,r,i,z,y$) obtained from the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS1; Chambers et al., 2016) or PS1 data archive, which is not available for the most southern galaxies. Additional photometry was obtained from the VISTA Hemisphere Survey (VHS; McMahon 2012), and the UK Infra-Red Telescope (UKIRT) Hemisphere Survey (UHS; Dye et al. 2018; Irwin 2013), with a limited, however, coverage of our sample for each survey [^5] For only one case, NGC 5253, due to the lack of PS1 data and the availability of VHS we used the y-VHS instead..
2.3 Removing foreground stars
At the start of the project, Gaia DR2 was available. This release included a meaningful number of measurements (for parallax and proper motion) only for a few galaxies. Therefore, we initially performed the foreground selection based on Gaia data for the galaxies M31 and M33, to derive specific cuts on parallax and proper motion (see Paper I). With the release of DR3 (Gaia Collaboration et al., 2023) both the number of available data increased significantly and the quality of the measurements has improved. Therefore, we decided to revisit this process.
We performed foreground removal for each galaxy independently, defining parallax and proper motion limits based solely on sources within each galaxy. For this process we defined an appropriate ellipse around each galaxy (using an optical image of the galaxy from the Digitized Sky Survey (DSS)-red) that would enclose the majority of its sources. This border does not coincide (necessarily) with any other radial profile derived from other means (e.g. $D_{25}$ or Petrosian radius), but it does not affect our results, which are based on statistical treatment of the data. We next defined a search radius for Gaia sources around each galaxy by taking a radius about twice larger than the semi-major axis of the ellipse defined around the galaxy (in contrast to the box defined around M31 and M33 in Paper I - however, the change does not affect the end result). We cross-matched (using a 1″  search radius) the Gaia queried catalog around each galaxy with the corresponding catalog we had built from IR and optical photometry. Most of the galaxies have Gaia data, but this depends on the distance and the crowding for each galaxy (number of Gaia sources are presented in Table 2).
Our purpose is to optimally select the foreground sources within the galaxy so that we eliminate (minimize, in reality) contamination, as foreground and galactic sources cannot be distinguished from the original IR catalogs. To achieve this we followed the Gaia source selection from Lindegren et al. (2021). For this, we first calculate the median error of each of the quantities parallax_error, pmra_error, and pmdec_error, and we used this to replace all corresponding values that were smaller than their median ones (as these might be underestimates of the real error). We excluded all sources without Gaia data, and then we selected sources with the following criteria: astrometric_excess_noise¡1, pmra_error / pmdec_error ¡3, parallax_error¡1.5, and phot_g_mean_mag $\leq 20.7$ (updated for DR3). Following Paper I, we first derived the distributions of the quantities parallax/parallax_error [^6] In Paper I we applied a 0.03 mas offset (Lindegren et al., 2018). However, it seems that its value is smaller and possibly consistent with 0 (Groenewegen, 2021). In any case, the inclusion of this offset would lead to a subtle difference in the parallax/parallax_error cut, which has no real impact in our work, and hence we did not apply it., pmra/pmra_error, and pmdec/pmdec_error for all sources outside the ellipse defined for each galaxy. These foreground distributions were fit with a spline function. Then, we plotted the equivalent distributions for the sources within the ellipse, which consist of the foreground population and the true galactic one. The foreground contamination is estimated by the (previously defined) spline scaled according to the ratio of the density inside and outside the galaxy. For the galactic population, we assumed a Gaussian distribution. The total distribution is therefore fit by combining this Gaussian along with the foreground contamination. After obtaining the mean value and standard deviation of this Gaussian profile, we determined as foreground sources those sources with values larger than $3\sigma$. For the parallax/parallax_error we considered as foreground sources only those with values larger than $+3\sigma$ since negative values are not valid measurements. The selected values for each galaxy are presented in Table 2. We also kept all sources without Gaia data, as there is no way we can decide whether such a source is foreground or galactic. By doing this, we actually increased the contamination but we were certain that we did not exclude possible interesting sources.
This step above was successful for the majority of the galaxies. However, there were cases where either the number of sources within the galaxy was too low (e.g. 17 stars in NGC 3077) or the distribution was very sparse and the fit was not trustworthy (e.g. IC 10 with 376 sources within the galaxy). For all these cases (11 out of 26 galaxies, indicated with an ’*’ in their parallax and proper motion limits in Table 2) we used the criteria derived from M31, which is the most populated galaxy.
During the cross-matching of the original IR catalogs with other surveys (including the Gaia DR2 initially) all multiple matches were removed (typically less than $\sim 3\%$; Paper I). Upon inspection of the results we found that a very small fraction of IR sources matched the same Gaia DR3 source. These sources were present in the initial point-source catalogs with different IDs, but they had similar coordinates and magnitudes. We considered them as duplicates in the initial catalogs. Although they matched the same Gaia source, the different magnitudes in IR could lead to different classification results. To avoid confusion we opted to remove these sources altogether. Depending on the initial IR catalog this fraction was less than $0.1\%$ (e.g. in NGC 4736 and M83 from Khan 2017, or NGC 6822, NGC 300, M81 from Khan et al. 2015), to $\sim 0.3\%$ (e.g. in NGC 253 and NGC 55; Williams & Bonanos 2016) and up to $\sim 7\%$ (e.g. in IC 10, IC 1613, and WLM; Boyer et al. 2015). The final number of sources per galaxy that was analyzed in this work is presented in the last column of Table 2.
Table 2: Number of sources per photometric survey and foreground selection criteria per galaxy. The columns contain the initial number of Spitzer sources as well as their cross-matches with Pan-STARRS, Gaia , UHS, and VHS. Parallax and proper motion limits for foreground detection are provided, before the final number of sources after removing duplicate sources.
Galaxy Spitzer PS-DR1 Gaia-DR3 UHS VHS Parallax/error pmRA/error pmDec/error Selected Final
WLM 14234 3331 399 0 3291 --0.08+0.99 0.26$\pm$1.20 --0.10$\pm$1.11 14091 13139
NGC 55 8698 0 729 0 0 --0.06+1.13 0.02$\pm$1.20 0.06$\pm$1.49 8524 8496
IC 10 32901 4547 1931 6006 0 --0.16+1.12* 0.09$\pm$1.15* --0.13$\pm$1.27* 31673 29499
M31 815811 410634 26332 387613 0 --0.16+1.12 0.09$\pm$1.15 --0.13$\pm$1.27 809142 809052
NGC 247 13398 2470 621 0 0 0.01+0.85 0.21$\pm$0.98 --0.03$\pm$1.14 13095 13095
NGC 253 8734 1578 522 0 0 0.06+1.20 --0.00$\pm$1.07 0.02$\pm$1.31 8409 8381
NGC 300 20511 0 1400 0 11480 0.00+1.20 0.14$\pm$1.12 --0.12$\pm$1.23 20161 20153
IC 1613 28371 10364 1229 0 0 --0.18+0.96 0.11$\pm$1.03 --0.01$\pm$1.04 28245 26396
M33 73206 52455 11049 47594 0 --0.09+1.09 0.11$\pm$1.12 0.04$\pm$1.17 71847 71847
Phoenix Dwarf 10831 0 499 0 1212 0.01+0.87 0.35$\pm$0.95 --0.19$\pm$0.95 10703 10021
NGC 1313 6156 6156 481 6156 6156 0.27+1.36* 0.06$\pm$1.41* 0.45$\pm$1.14* 5970 5970
NGC 2366 495 156 64 0 0 --0.16+1.12* 0.09$\pm$1.15* --0.13$\pm$1.27* 462 462
NGC 2403 16644 3735 1517 0 0 --0.01+1.43 0.26$\pm$1.64 --0.05$\pm$1.47 15936 15910
M81 28479 3894 1072 0 0 --0.16+1.12* 0.09$\pm$1.15* --0.13$\pm$1.27* 27895 27875
Sextans B 4914 1166 141 0 0 --0.16+1.12* 0.09$\pm$1.15* --0.13$\pm$1.27* 4852 4413
NGC 3109 9474 2988 1069 0 0 --0.09+0.93 --0.01$\pm$1.06 --0.05$\pm$1.18 8939 8935
NGC 3077 2617 271 90 0 0 --0.16+1.12* 0.09$\pm$1.15* --0.13$\pm$1.27* 2548 2548
Sextans A 2888 880 219 0 355 0.03+0.84 --0.20$\pm$0.98 --0.12$\pm$0.94 2848 2693
NGC 4214 1159 368 95 89 0 --0.16+1.12* 0.09$\pm$1.15* --0.13$\pm$1.27* 1149 1149
NGC 4736 10043 1248 349 657 0 --0.16+1.12* 0.09$\pm$1.15* --0.13$\pm$1.27* 9861 9861
NGC 4826 4659 480 149 306 0 --0.16+1.12* 0.09$\pm$1.15* --0.13$\pm$1.27* 4577 4575
M 83 15020 2422 1396 0 3877 --0.04+0.76 0.31$\pm$1.20 0.07$\pm$1.05 14132 14132
NGC 5253 721 0 119 0 187 --0.16+1.12* 0.09$\pm$1.15* --0.13$\pm$1.27* 622 622
NGC 6822 25599 18659 7061 0 15205 0.01+1.13 --0.08$\pm$1.05 --0.24$\pm$1.16 22483 22471
Pegasus DIG 11316 2234 251 0 0 --0.16+1.12* 0.09$\pm$1.15* --0.13$\pm$1.27* 11147 10530
NGC 7793 5535 887 353 0 0 0.04+1.47 0.09$\pm$1.32 -0.04$\pm$1.06 5433 5425
777 Spitzer data for galaxies IC 10, IC 1613, Pegasus DIG, Phoenix Dwarf, Sextans A and B, and WLM are derived from Boyer et al. (2015); M33, M81, NGC 2403, NGC 247, NGC 300, NGC 6822, and NGC 7793 from Khan et al. (2015); NGC 2366, NGC 253, NGC 4214, NGC 5253, and NGC 55 from Williams & Bonanos (2016); M31, M81, NGC 3077, NGC 4736, NGC 1313, and NGC 4826 from Khan (2017). \$*\$\$*\$footnotetext: For those galaxies the parallax and proper motion criteria derived from M31 were used, as there were not enough data to determine these quantities by using their data only.
2.4 Quality cuts
The application of the classifier was straightforward. After building a source catalog for each of our galaxies we parsed it to the classifier, which identified and built the necessary features, i.e. calculated the color terms $r-i$, $i-z$, $z-y$, $y-[3.6]$, $[3.6]-[4.5]$. We used the pre-trained models from Paper I to obtain the predictions for each galaxy [^8] For the application of the classifier check the GitHub page of the project (https://github.com/gmaravel/pc4mas), where a python notebook and the models can be found..
The classifier provides a classification among the following classes (see Paper I for details): BSG - a rather loose class of early-type and hot stars, YSG - evolved yellow type stars, RSG - red supergiants, BeBR - B[e] supergiants, LBV - Luminous Blue Variables, WR - Wolf-Rayet stars, GAL - galaxies including AGNs and background QSOs (i.e. all extragalactic outliers). The probability per class is provided for each of the algorithms used (i.e. Support Vector Machines, Random Forest, Multi-Layer Perceptron; see Paper I for more details), which are then combined (with equal weighting) to provide the final set of probabilities. For each source the final (predicted) class corresponds to the one with the highest probability of the ensemble approach (see Paper I for details).
In Paper I we investigated the number of correct vs. incorrect sources with probability. We found that the mean values for the correct and incorrect classifications (in the combined sample of M31 and M33) were $0.86\pm 0.01$ and $0.60\pm 0.03$, respectively. A cut at 0.86 would significantly limit the sources to be considered since the classifier performs slightly worse for lower metallicities than M31 and M33. Therefore, we opted to use a cut at 0.66, equivalent to the mean value $+3\sigma$ of the incorrect classification distribution, which gave us the opportunity to consider a larger sample of sources while keeping the fraction of incorrectly classified sources relatively low (c.f. Fig. 6 and 8 of Paper I).
Additionally, band availability (i.e. missing data) significantly affects the performance and robustness of the prediction model, specifically the probability. Our tests on missing data imputation (c.f. Fig. 9 in Paper I) indicate that the classifier can effectively manage up to two missing features (equivalent to three missing bands, since each feature is a color index) with a loss of accuracy of approximately 10%. This resilience extends to three missing features, albeit with an increased accuracy loss of less than 20%. More missing values lead to an approximate 40% loss, hence making the classifier unreliable.
Given that missing values are a common occurrence for our sources, we identified two scenarios for selecting objects for further processing. The more balanced approach involved selecting sources with a final probability greater than 0.66 and band completeness exceeding 0.6 (i.e. missing two features, three bands in total). Alternatively, a more relaxed option considered sources with a probability greater than 0.50 and band completeness exceeding 0.4 (i.e. missing three features and 4 bands). However, the second scenario introduces a higher rate of false positives, increasing noise, due to more misclassifications, in the final results. Therefore, we decided to proceed exclusively with the first approach. To ensure flexibility, our published catalogs include data for all available sources, allowing users to apply their own selection criteria as needed.
2.5 Collecting known sources from literature
To better understand the strengths and the limitations of our classifier we needed to compare its predictions with previously classified sources. For this reason, we undertook the very demanding task to search the literature for all known sources found in our galaxy sample. We limited our search to sources with secured spectral classification, i.e. that had been obtained with spectroscopy and not from photometric or other (e.g. variability) criteria. We have collected the complete spectral samples - to the best of our knowledge - for all the 26 target galaxies, accounting to 5273 sources (from 83 different works). This number includes all massive stars known, including candidates, as well as another $\sim 330$ point sources (such as carbon stars, background galaxies, Hii regions, planetary nebulae, clusters). The numbers and corresponding references per galaxy can be found in Table 7 of the appendix. We note here that no spectral classifications were available for 5 galaxies (i.e. NGC 2366, NGC 3077, NGC 4214, NGC 4826, NGC 5253). In all cases we carefully checked for and removed duplicates, keeping the most recent and precise classifications. The current catalog serves as the most complete source of reference for spectral classifications for massive stars and candidates.
3 Results
In this section, we provide the results of the application of the machine-learning classifier to our galaxy sample, in the form of a full catalog, a table providing the statistics per class (per galaxy) based on the most secure predictions, as well as some indicative color-magnitude diagrams.
3.1 Catalog description
In Table 3 we provide the (first few lines) of the final catalog for all sources from all galaxies, comprising 1,147,650 sources and spanning 78 columns. In the Table we provide the source ID, Spitzer coordinates, Gaia DR3 ID, proper motion and parallax, Spitzer , PS1, and near-IR photometry and errors, previous classification (if available), probabilities per class for each method including the ensemble one, as well as final class, final probability, and band completeness.
Table 3: Final source catalog with predicted classifications for all galaxies.
ID R.A. (J2000) Dec. (J2000) Gaia_DR3_ID Final_Class Final_Prob Band_Compl
(deg) (deg)
WLM-2 0.52017 -15.44622 WR 0.602 1.0
WLM-3 0.52012 -15.40681 WR 0.647 0.4
WLM-4 0.52012 -15.36033 RSG 0.859 1.0
WLM-5 0.52012 -15.46578 WR 0.618 0.2
WLM-7 0.52012 -15.44331 WR 0.648 0.2
WLM-8 0.52012 -15.41617 WR 0.574 0.2
WLM-9 0.52008 -15.50714 WR 0.612 0.2
WLM-10 0.52008 -15.46256 WR 0.536 0.2
WLM-11 0.52008 -15.56622 WR 0.438 0.6
WLM-12 0.52008 -15.42581 WR 0.615 0.2
WLM-13 0.52008 -15.43997 WR 0.613 0.2
WLM-14 0.52008 -15.54617 WR 0.558 0.2
WLM-16 0.52008 -15.56514 RSG 0.396 0.2
WLM-17 0.52004 -15.53664 YSG 0.476 0.6
WLM-18 0.52004 -15.55456 WR 0.473 0.2
WLM-19 $0$15.40131 WR 0.430 0.6
WLM-21 $0$15.46006 WR 0.646 0.2
WLM-22 $0$15.60042 WR 0.615 0.2
WLM-24 0.51996 -15.45511 WR 0.616 0.2
WLM-25 0.51996 -15.35928 RSG 0.487 0.2
 [^9] This table is available in its entirety in the VizieR/CDS catalog tool. A portion is shown here for guidance regarding its form and content.
3.2 Populations
Table 4: Number of predictions per galaxy, after applying the selection criteria on probability and band completeness (from Sect. 2.4).
Galaxy Total RSG YSG BSG BeBR WR LBV GAL
WLM 526 268 85 13 2 147 1 10
IC 10 1622 658 11 0 0 947 0 6
M31 225176 81734 696 268 33 140153 12 2280
NGC 247 897 372 17 6 0 423 1 78
NGC 253 385 167 36 1 1 118 5 57
IC 1613 2964 2351 392 39 1 162 5 14
M33 31635 25808 322 212 23 4767 8 495
NGC 2366 42 29 5 0 0 2 0 6
NGC 2403 950 620 74 10 1 217 6 22
M81 1387 382 36 1 2 899 1 66
Sextans B 231 176 25 2 0 24 0 4
NGC 3109 736 363 38 7 0 319 1 8
NGC 3077 96 17 1 0 0 68 0 10
Sextans A 168 100 20 8 0 35 2 3
NGC 4214 142 42 10 4 0 81 0 5
NGC 4736 341 49 30 0 1 227 0 34
NGC 4826 151 24 19 0 0 94 0 14
M83 565 104 28 10 4 388 1 30
NGC 6822 8007 6992 220 34 4 690 17 50
Pegasus DIG 627 217 15 1 0 389 0 5
NGC 7793 9 6 2 0 0 1 0 0
TOTAL 276657 120479 2082 616 72 150151 60 3197
For further consideration and exploration of the results we only kept sources that satisfied the quality criteria as defined in Section 2.4. This allowed us to exclude the most uncertain predictions from the classifier, either due to the low probability or the number of missing values (below the band completeness threshold we set). These limits remove the noise and allow us to investigate the results and assess the performance of the classifier. Table 4 presents the total number of source classifications, along with the number of predictions for the classes of RSG, YSG, BSG, BeBR, WR, LBV, and GAL, independently.
Out of the 26 galaxies, 5 of them (namely NGC 55, NGC 300, Phoenix Dwarf, NGC 1313, NGC 5253) have zero sources (hence are not presented in Table 4). This is due to a lack of PS1 data which leads to more missing values than what we accepted (less than 0.6). Since only the IR data is available for these galaxies, the predictions would be highly uncertain and - therefore - were excluded from further analysis.
Overall, we find the highest number of predictions to be for RSGs and WR stars. In particular, M31, M33, NGC 6822, IC 1613, IC10, and M81 have more than 1000 classified sources. However, we already know (from Paper I) that WR predictions suffer from many false positives. We further discuss the performance of the classifier in Section 4.1.
3.3 Color-Magnitude Diagrams
Figure 2: Example of CMDs ($z$ vs. $r-z$ and [4.5] vs. $[3.6]-[4.5]$) for M31 (top) and NGC 2403 (bottom) with predicted classifications. We plot all sources that satisfy the quality cuts imposed in Section 2.4, and their photometric values come from the original data (i.e. we do not plot sources whose values have been imputed during the application of the classifier). We use cyan circles for Blue Supergiants (BSG), blue diamonds for B[e] Supergiants (BeBR), gray triangles for galaxies (GAL), green pentagons for Luminous Blue Variables (LBV), red crosses for Red Supergiants (RSG), purple empty triangles for Wolf-Rayet stars (WR), and yellow filled bottom-sided for Yellow Supergiants (YSG). The total number of sources per class is provided next to the class in the legend.
We explored the locations of the predicted populations in Color-Magnitude Diagrams (CMDs). Fig. 2 shows optical and mid-IR CMDs for two example galaxies, M31 (as the most populated galaxy) and NGC 2403. We note that we do not plot sources whose values have been imputed during the application of the classifier, but only sources with their original photometry. We use different symbols to indicate the predicted class for each source, and we show the total number of predictions per class in the legend.
We find that the position of sources matches their predictions. In the $z$ vs. $r-z$ CMD we see BSG located on the left (of approximately $r-z\sim 0$) while YSG and RSG extend to redder colors as expected. The few BeBR and LBV predicted are located (consistently) close to the BSG and WR. The latter ones display a much broader distribution which is not real in many cases, as this class suffers the most from false positives. This is more striking in the case of M31, where we note the ”line” of WR that extends from the bulk of sources (with $r-z\sim 0$ mag) to a point of $r-z\sim-5$ mag (and of $z\sim 27.5$ mag). Most probably the classifier confuses these faint but blue sources with WR, because M31 is a nearby galaxy and we can observe fainter stellar populations. In contrast, we do not see this in NGC 2403.
Regarding the [4.5] vs. $[3.6]-[4.5]$ CMD we notice that the majority of the RSGs lie around $[3.6]-[4.5]\sim 0$ mag in both M31 and NGC 2403, consistent with what we would expect. For M31 though there are sources with bluer colors. Although there are indeed RSGs with bluer colors (e.g. Bonanos et al. 2009; de Wit et al. 2024) there is confusion with other populations (due to the proximity of M31). There is definitely confusion of WRs with other populations (WR is the class with the higher false positive results). In NGC 2403 WRs extend to slightly redder colors which is consistent with what we would expect. YSGs are located around 0.0 for both galaxies. However, the two very bright points in NGC 2403 are most probably foreground sources. In this CMD, GAL seems to become more evident from the bulk of the stellar populations and they are distinctly separated with $[3.6]-[4.5]>0.5$ mag.
Overall, the position of the sources matches their predicted classification. There is of course a fraction of sources that is misclassified. This is due to a number of physical and technical reasons, such as the fact that strict boundaries in these parameter spaces between some classes do not exist (e.g. BSG and WR, YSG, and RSG), the fact that there could still be some contaminant sources either because these have not been part of the classifier’s training, or because there could be photometric errors (e.g. crowding, aperture contamination, or remaining foreground sources). We further explore these in Section 4.1.
4 Discussion
In this section, we compare our predictions with literature data to understand the performance of the classifier, explore the trends of the various classes with metallicity, and discuss the luminous RSGs and dusty YSGs in our sample.
4.1 Performance: comparison to literature
Figure 3: Success rate vs. distance (top panels) and metallicity (bottom panels), using a uniform prior (left panels) and a unimodal beta distribution (right panels) with a peak corresponding to $77\pm 7$% (based on the performance of the classifier during developing). We notice a small decrease of the success rate with distance and a relatively flat behavior with metallicity, especially in the case where a prior is implemented (see Section 4.1 for details). The number of available classified sources from the literature is indicated by the size of the points and the colorbar on the right.
To properly assess the performance of our classifier when comparing it with literature results (which are taken as ground truth) we considered two factors. The first was to provide a comparison by selecting the best candidates as defined from the quality criteria in Section 2.4 (i.e. final probability of ¿ 0.66 and band completeness ¿ 0.6). The second was to estimate the classification error, based on the probability distribution for each source.
Since the classifier provides a probability for each of the classes, the probability of two (or more) classes may be similar. In that case, although a single probability (and therefore class) is returned by the classifier, it does not mean that another class(es) is not possible (e.g. consider the case where we have 0.45 and 0.40 for a RSG and YSG class, respectively). To account for this we decided to estimate the class error by keeping all possible classes for which the probability did not differ significantly. For this, we first sorted all classes based on their probability (from highest to lowest). The first class in this sorting was always the one to be returned. If the ratio of the probability of the second to the first class differed more than a factor of two, it was rejected, and the first class was the only result. If the ratio of probabilities is less than two, the second class was added to the list of possible classes. We continued this process up to the point where the next class was rejected. Thus, we created a list of possible classes. When comparing our predictions to literature classes we identified the following cases:
First, an exact match when the first class of the list was the same or equivalent to the literature class. That was the case for straightforward classifications such as those used to train the classifier (see Paper I, table 2 and Section 3.2). For example, an A5 Ib or B9 Ia star predicted as a ’BSG’ would be an exact match. Similarly, a K2/3 I star or a carbon star predicted as a ’RSG’. Although the last pair (RSG and carbon star) is not a physical exact match, the classifier has been trained on a feature scheme (color indexes) for which these two classes overlap. In other words, due to the lack of features and data, the classifier cannot separate these two classes. Another example is when a rather broad classification is provided by the literature, such as ”emission line star”, which can be consistent with any of the classes of ’LBV’, ’BeBR’, ’WR’, and ’BSG’ [^10] This information was cross-matched with the source papers to verify their stellar nature. For example, the classification of ’EmObj’ from Bonanos et al. (2024) refers to objects that although display emission lines they are of uncertain nature. Therefore, these cases were considered uncertain.. For our classification scheme, these are equivalent cases and therefore correct whenever present.
Second, a match within the error when any class for the list of possible results was similar or equivalent to the literature class. That is the case when, for example, for a literature class of A5 II we had a predicted list of ’LBV’, ’YSG’, ’BSG’, or ’WR’. Although the first result ’LBV’ (which was the one returned as the final class) did not match the literature, there was a class within the accepted ones (i.e. ’BSG’) that matched the literature. Therefore, our prediction was correct within its error.
Third, there were cases where the literature result was too uncertain to safely match any of our classes. Such an example is the ’Composite’ classification assigned to the source B17 in WLM (Britavskiy et al., 2015), which does not provide any further information on how these sources can be compared to our classes.
Fourth, cases where none of the predicted classes matched the literature one. These are obvious cases such as a ’late G’ star or a carbon star predicted as ’WR’.
Fifth, a no-match result was returned when there was no match within 1″  between the literature source and our catalog. These sources were excluded from further analysis.
We note here that when performing this comparison of spectral classification and the prediction from our photometric classifier, variability may introduce an offset. Especially for cases of sources with significant variability (e.g. LBVs, RSGs) the epochs of spectroscopic and photometric observations (which are not concurrent in the optical and IR) may lead to a mismatch between the spectral type and the prediction. For larger populations though (such as RSGs) we expect statistically consistent results.
In Appendix B we provide in detail the statistics of matched sources per galaxy. In Fig. 3 we plot the success rate (corresponding to the number of correct and correct within error) over the subsample of literature sources, which matched best candidates from our catalog with distance and metallicity of each galaxy. We have taken extra care to deal with the errors in these fractions for two reasons. First, the fractions are bounded (0 to 1) so their error cannot extend beyond these values. Standard approaches of error estimate on the fractions (such as Wald or Wilson methods) would yield symmetric errors, leading to values exceeding the bounds (e.g. close to 1). Secondly, the sample size is different in each galaxy. Based on the current sample size and the number of correct predictions we get only one result (of the possible combinations) for the success rate, which follows a binomial distribution. To overcome all these and to avoid any Gaussian approximation, we have followed a Bayesian approach (see Appendix C for a more technical overview). We estimate the posterior probability of the success rate based on the likelihood provided by the binomial distribution (given the sample size and the correct predictions in each case) and a prior. The prior corresponds to the knowledge we have regarding the distribution of probable success rates for our classifier. We considered two options for the prior, a uniform one (i.e. single success rate value), and a unimodal beta distribution (plots in Fig. 3 left and right correspondingly; only galaxies with literature data and good candidates are plotted). This option translates to a success rate that has a single peak and drops off towards both bounds. We already have data for this from Paper I, in which we have concluded that the accuracy was 0.83 for the M31 and M33 galaxies (on which the classifier was trained) and 0.70 for the test galaxies (i.e. IC 1613, WLM, and Sextans A). We can therefore construct a distribution with a mode at a mean value of 0.77 and a variance of 0.01111111The difference between the two sample values is $\sim 0.07\%$, which corresponds to a variance of 0.005. As this value is indicative of how the distribution drops from the peak value we opted to relax it, and double the value to 0.01, to allow for a smoother distribution to include more probable success rates.. Given the likelihood and the prior we can construct the posterior probability distribution per galaxy. From this, we can extract basic statistical properties including asymmetric errors corresponding to the highest posterior density interval set at 95%.
By examining the left panels of Fig. 3 we first notice that the success rate of the classifier drops with distance, although not significantly. This is expected, as the farther the galaxy is, the greater the level of confusion. The most extreme example is M83 for which we have,additionally, a small number of good candidates (fulfilling the criteria of Section 2.4). Interestingly, we also note that the classifier seems to be rather robust with metallicity. We achieve good results even at the lowest metallicity environments. There is an exception of three galaxies (namely NGC 7793, NGC 2403, and NGC 253) for which the success rate is lower than 0.4 (when using the uniform prior). These are the galaxies with the smallest number of classified samples for which we got very low success rates (e.g. one correct prediction out of the 6 sources in NGC 253, 0 correct out of 2 in NGC 7793, and 2 out of 7 in NGC 2403). These cases are also sensitive to the low-number statistics since a small change can lead to a significant change in the success rate. For example, in NGC 253 there were a couple of sources which were found uncertain (such as ’cluster’ or ’nebula’, which the classifier cannot predict). If we were to remove these two from the sample, then the original success rate would change from one out of 6 to one out of 4 sources (from 0.16 to 0.25), which is a significant change. On the other hand, Sextans B has only two classified sources that are both predicted correctly. Hence we ended up with a 100% success rate which was just a random realization of the binomial distribution.
By incorporating the prior from the unimodal beta function that peaks at a mean success rate (as defined previously) we notice (see right panel of Fig. 3) that the values are updated and both extreme values (i.e. NGC 7793, NGC 253, NGC 2403, and Sextans A) are driven towards to more probable (and realistic) values, within the 60–94% success rate. The small decrease of the success rate with distance is still evident, but the performance of the classifier with respect to the metalliciy flattens. M31 and M33 stand out in these plots because they both have the largest numbers of sources by far ($\sim 430$ and $\sim 380$, respectively, resulting in small errors), and the best success rates. The latter is to be expected since the classifier has been trained on samples from these two galaxies.
In summary, we conclude that our classifier performs exceptionally well even at lower metallicities (e.g. $0.07\,Z_{\odot}$ for Sextans A and Phoenix Dwarf), despite not being explicitly trained for them. There seems to be a greater dependence on distance (due to confusion) which makes it more efficient at distances smaller than 1.5 Mpc but with only a minor loss beyond 3 Mpc (where the furthest galaxies are located).
4.2 Populations with metallicity
Figure 4: The fraction of predicted population with metallicity per class. Despite the presence of large errors and important physical and observational biases, there are noticeable trends of each population with metallicity (see Section 4.2 for details). The symbol size and color for each galaxy reflects the corresponding sample size. For clarity we assign each galaxy to an integer ID, shown in the legend at the top.
In this section, we examine the trend of each class separately with metallicity. But before presenting our results, we describe in detail our biases.
We repeat here that our source selection was based on Spitzer point-source catalogs. That means that the catalog comprises of sources visible in the IR, which means that a certain fraction of some populations (e.g. OB main-sequence stars, stellar sources without dusty environments) are not included. Given this, we understand that the BSG class, for example, does not reflect the complete sample for any galaxy. That is evident if we compare the most recent catalog of OB stars in Sextans A by Lorenzo et al. (2022) with ours. We find matches for $\sim 28\%$ (of 106 sources in total, within a 1″  search radius). Another factor, influencing the fraction of populations we can observe, is the galaxy inclination. For example, NGC 55 is almost edge-on which means that a (significant possibly) fraction is not visible. While in other cases, such as in Sextans A which is a smaller and less crowded galaxy, we can pinpoint all sources. The difference between these two galaxies highlights that the galaxy’s type and star formation history have an important impact also. A recent star-formation event will lead to the birth of new stars which would (most probably and depending on the metallicity) lead to increased numbers of WR stars. As time passes these populations (along with OB main sequence stars) will decrease, and in turn the populations of more evolved phases such as YSG and RSG will increase.
Following the above physical reasons there are several observational biases. One main constraint is the limitation in the pointings of Spitzer and its coverage, which affects the completeness. In Maravelias et al. (2023) we noted that in NGC 55 there are four known LBVs and our approach was successful to recover two of them. The other two were not recovered due to their location, which lay outside the observed Spitzer fields. Therefore, an unknown fraction of all populations are affected. This is more critical in the rarest cases (such as for LBVs and BeBRs) while probably not significant for the most populous ones (such as RSGs, and BSGs). Our classifier is based on the presence of photometry in certain Spitzer and Pan-STARRS bands. Additionally, many sources do not have photometry in all these bands, and, consequently, the quality of their predictions did not pass our quality criteria. Although this compensates for not adding erroneous predictions, there is an important fraction of sources not considered in our analysis.
Considering all the above caveats, Fig. 4 presents the fraction of predictions per class over the total sample of only the best candidates. Our results are provided using the same approach for error determination described in the previous section (using the 95% interval), but we are using a separate prior for each class. The peak and the variance of the distribution for each class are provided from the results we have obtained in Paper I (c.f. table 5) [^12] For the LBV class we used the fractions and corresponding variance derived from the Support Vector Machine methods, the only non-zero result.. We removed galaxies without any good matches (NGC 55, NGC 300, Phoenix Dwarf, NGC 1313, NGC 5253).
Starting with LBVs we cannot draw secure conclusions regarding their trend with metallicity, given their small numbers and large errors. There is some indication that lower-Z environments may host more LBVs (e.g. Sextans A).
The B[e] supergiants (BeBR) is another rare class of objects, for which there are not many confirmed cases in low metallicity galaxies (see e.g. in Maravelias et al. 2023). These sources are (by definition) very bright in IR due to the dust formation in their complex circumstellar rings, which means that we should be able to recover almost their total populations across all galaxies. Therefore, the slight increase with decreasing metallicity should be real. Whether or not we expect a larger number of such objects in low metallicity environments is currently unknown (because of the lack of their formation channels; see e.g. Kraus 2019).
Metallicity has a significant impact on stellar winds of hot stars, which are mainly driven by iron. Therefore, decreasing metallicity leads to less efficient winds, resulting in fewer stars managing to expel their outer layers to become WR. This is depicted in the plot by the decrease of the WR fraction with metallicity, which is slightly more evident below 0.4 Z. Given that the fraction of false positives within WR is quite high it is possible that their real fraction is even smaller than what depicted in the plot. Moreover, as they correspond to young massive stars their populations correlate also with regions of recent star formation events (such as starburst NGC 4214; Williams et al. 2011).
As the metallicity decreases, stars lose their angular momentum much slower, preserving in this way their initial rotational speeds longer and increasing the internal mixing (the stars become more compact and have higher temperature and density in their centers). Consequently, there seems to be an extension of the main-sequence phase (hydrogen burning at the core; Georgy et al. 2013; Ekström et al. 2012). Simultaneously, it is possible to have He burning happening much earlier (i.e. higher temperatures), before the RSG phase (Yoon et al., 2008). Those stars will spend more time as YSGs rather than RSGs. Moreover, the more massive and luminous RSGs have strong winds that lead to a fast stripping of their envelopes, resulting in post-RSGs looking like BSG or YSG (Massey et al., 2023; Zapartas et al., 2024). Ultimately, these would lead to more BSGs and YSGs present at any time, than their corresponding numbers in higher metallicity galaxies. This is indeed observed in Fig. 4.
The RSG population is not affected significantly (if any) by the metallicity (e.g. Antoniadis et al. 2024, 2025). Therefore, we would expect a rather flat trend or slight drop with metallicity. However, this is not what we see in the RSG panel of Fig. 4. We attribute this to an observational bias, as we constructed our catalog based on Spitzer sources corresponding mainly to evolved and dusty sources. At lower metallicity galaxies, more stars spend their majority of their time in the BSG/YSG phase, which means that the number of these sources with dust will drop. Consequently, the ratios of RGSs over the total number of IR sources (which includes fewer actually observed BSG/YSG as the metallicity decreases) lead in an upward trend.
These intriguing findings reveal the impact of metallicity on massive star populations. They accentuate the need for more comprehensive studies to overcome current limitations imposed by the physical and observational biases of our approach.
4.3 RSG luminosity functions
Figure 5: Luminosity functions for all sources identified as RSGs in M31, M33 and NGC 6822. We note the presence of some very luminous sources for M31 and M33 with $\textrm{log}(L/L_{\odot})\,$¿5.5, indicated by a dashed line (see Section 4.3).
Table 5: Properties of luminous predicted RSGs in M31 and M33.
ID Gaia_DR3_id Parallax pmRA pmDec Origi-l Fi-l Fi-l Band $\textrm{log}(L/L_{\odot})\,$
Class Class probability completeness
(mas) (mas yr-1) (mas yr-1) (dex)
M31-439614 RSG 0.94 0.6 6.8
M31-439351 RSG 0.67 1.0 6.8
M31-439254 RSG 0.79 1.0 6.7
M31-350 381305998254302336 --0.056 0.21 0.089 M1I RSG 0.97 1.0 5.8
M31-439595 381275589889380352 RSG 0.87 1.0 5.8
M31-651 RSG 0.97 1.0 5.7
M31-569 387318505783297792 RSG 0.98 1.0 5.7
M31-690 381285378106567680 RSG 0.96 1.0 5.7
M31-439785 381168662371881088 --0.24 0.55 --0.5 RSG 0.87 1.0 5.7
M31-519 381317508755717888 RSG 0.98 1.0 5.7
M31-440141 381172231487764352 RSG 0.98 1.0 5.7
M31-688 375276104678171392 RSG 0.98 1.0 5.7
M31-787 381285279321284352 RSG 0.9 1.0 5.6
M31-439741 381171067553249920 --0.034 0.013 0.04 RSG 0.91 1.0 5.6
M31-746 375296067688047616 RSG 0.70 1.0 5.5
M31-440112 381133306201068416 0.03 --0.63 --1.5 RSG 0.76 1.0 5.5
M33-179 303288535788135680 RSG 0.84 1.0 5.8
M33-520 303267473268558208 RSG 0.98 1.0 5.6
M33-173 303365948277821952 --0.18 0.099 --0.18 OB RSG 0.96 1.0 5.5
M33-646 315403264139899008 RSG 0.98 1.0 5.5
M33-249 303283553626082944 --0.076 0.12 0.072 RSG RSG 0.93 1.0 5.5
M33-208 303378072972636032 --0.062 --0.16 --0.023 RSG RSG 0.98 1.0 5.5
Given that the highest success rate is obtained for RSGs we opted to explore their luminosity functions for the galaxies with sufficient numbers, namely M31, M33, and NGC 6822. Although historically the Humphreys-Davidson limit (Humphreys & Davidson, 1979) is found at $\textrm{log}(L/L_{\odot})\,$$\sim 5.8$ dex, more recent works lowered the limit to $\textrm{log}(L/L_{\odot})\,$$\sim 5.5$ dex (Davies et al. 2018 for the LMC and SMC; McDonald et al. 2022 for M31). Hence, it is very intriguing to investigate predictions of RSGs above this limit.
For this, we need first to estimate the luminosity for each source. We employed the bolometric correction for RSGs from Neugent et al. (2020) following their section 5, given the $K_{s}-\textrm{magnitude}$ as provided by the UHS survey (using a 3” aperture for $JK_{s}$ photometry). To correct for extinction, we assumed a uniform value of $A_{V}=0.75$ mag as they did, but for all sources. When not applying additional extinction for the brighter sources, our estimated luminosities are similar to those from (Wang et al., 2021), which are calculated by integrating the spectral energy distributions, a more robust method that takes into account the emission from the dust shell. For NGC 6822, we assumed $E(B-V)=0.25$ mag (Massey et al., 2007b). The luminosities of RSGs in NGC 6822 are underestimated by 0.05 dex compared to the luminosities of the RSGs in common with Antoniadis et al. (2025), which were calculated by integrating the spectral energy distributions. The results are presented in Fig. 5 (we note here that the uncertainty in this estimate due to the correction is 0.05 dex; Neugent et al. 2020).
From this plot, it is striking that there are several luminous sources above $5.5\,L_{\odot}$. In particular we find 22 M31 and 6 M33 sources. In Table 5 we present them by including their Spitzer and Gaia DR3 IDs, their Gaia parallax and proper motions (whenever available), previous (original) and final (predicted) classifications, as well as their final probability, band completeness, and estimated luminosity.
In M31, the three most luminous sources (close to $\textrm{log}(L/L_{\odot})\,$$\sim 7$; IDs: M31-439614, M31-439351, and M31-439254) do not have a Gaia counterpart, which means that these could be foreground red stars. However, the next most luminous source (ID M31-350) is a confirmed M1 I RSG (J004428.48+415130.9), with a range of $\textrm{log}(L/L_{\odot})\,$$=5.43-5.64$ McDonald et al. (2022). Such luminous sources may actually host a dense and complex circumstellar environment (similar to what has been observed in WOH G64; Ohnaka et al. 2024; Munoz-Sanchez et al. 2024a) that can contribute up to 0.3 dex in luminosity. If we consider these uncertainties in our approach, then our estimate of $\textrm{log}(L/L_{\odot})\,$$=5.9$ could be corrected by 0.35 dex, making it consistent with values from the literature. Based on our Gaia foreground cleaning method, the parallax and proper motion measurements make it a strong M31 candidate. Therefore, it certainly provides an upper bound for the Humphreys-Davidson limit in M31. Below that there are an additional 12 sources (between $\textrm{log}(L/L_{\odot})\,$$\sim 5.8-5.5$), of which 3 have Gaia parallax and proper motion values (making them best candidates of high luminous RSGs [^13] Among these sources is J003951.33+405303.7 from McDonald et al. (2022), corresponding to M31-439741, which we predicted robustly as RSG. Sources J004731.12+422749.1 (M31-1011) and J004539.99+415404.1 (M31-336) were also predicted as RSG, but we were unable to estimate their luminosity. For sources J004428.12+415502.9 and J004520.67+414717.3 we were unable to find a good match within a few arcsec in our catalog, probably due to our different Gaia cleaning approach., another set of 8 sources with Gaia photometry only (with at least one band measurement), and one source without any Gaia match (hence, more probable to be a foreground source).
In M33, we find two sources (M33-179 and M33-520) at $\textrm{log}(L/L_{\odot})\,$$\sim 6$). Since both of them lack Gaia parallax and proper motion measurements (but have photometric measurements) we cannot determine if they are genuine M33 members or not. These are followed by source M33-173 at $\textrm{log}(L/L_{\odot})\,$$\sim 5.7$, for which full Gaia information exists. However, this is a known OB star (Massey et al., 2016) and marks an erroneous prediction as a RSG. Another source (ID M33-646) has only Gaia photometry and, therefore, we can only tentatively consider it as a M33 RSG. The remaining two sources have full Gaia information, and have been classified spectroscopically as RSGs. Therefore, they are the sources with the most robust data regarding their M31 membership status and spectral classification, and their derived luminosity sets a limit at $\textrm{log}(L/L_{\odot})\,$$\sim 6$ (which could be lowered to $\textrm{log}(L/L_{\odot})\,$$\sim 5.7$ if we consider the correction due to the circumstellar contribution).
This exploration emphasizes the need for further detailed investigations into these luminous RSGs, to confirm their true nature as RSGs, and to accurately determine their luminosities through spectral energy distribution fitting. Precise luminosity measurements will also aid in exploring the Humphreys-Davidson limit (Humphreys & Davidson, 1979). JWST is the perfect instrument to provide images that will decisively determine (if and) how much these sources are contaminated.
4.4 Dusty YSGs
Figure 6: IR color ($[3.6]-[4.5]$) distribution of all (best candidate) sources predicted as YSGs. The violet curve corresponds to a Gaussian fit and the black dashed line indicates the $1\sigma$ threshold (0.24 mag) above which we identify sources as dusty YSGs.
Table 6: IDs of sources predicted as YSGs and identified as dusty.
ID ID ID ID ID ID
SextansA-2751 NGC2403-11707 M31-315890 IC1613-2519 IC1613-27853 NGC6822-19461
M81-12028 M83-5253 M31-318902 IC1613-3532 IC1613-27906 NGC6822-19716
NGC253-1672 M83-7102 M31-373976 IC1613-4592 IC1613-28726 NGC6822-20136
NGC253-1873 M83-13705 M31-389826 IC1613-5907 IC1613-29826 NGC6822-20875
NGC253-2299 M83-22806 M31-408762 IC1613-6759 IC1613-30616 NGC6822-20905
NGC253-3168 Pegasus-11040 M31-439225 IC1613-7221 IC1613-31585 NGC6822-21661
NGC253-3657 Pegasus-15252 M31-439242 IC1613-8299 WLM-89 NGC6822-21893
M33-57845 Pegasus-15778 M31-439244 IC1613-8407 WLM-234 NGC6822-22456
M33-58478 M31-5753 M31-445810 IC1613-8919 WLM-252 NGC6822-23887
M33-59527 M31-5947 M31-445852 IC1613-10537 WLM-13718 NGC6822-24378
M33-61917 M31-16636 M31-487362 IC1613-12631 WLM-16056 NGC6822-24663
M33-66957 M31-53947 M31-510271 IC1613-12668 WLM-19863 NGC6822-24791
M33-68837 M31-134734 M31-517843 IC1613-13160 NGC3109-3334 NGC6822-24814
M33-69046 M31-145674 M31-523223 IC1613-13595 NGC3109-4208 NGC6822-25657
M33-70621 M31-146048 M31-625493 IC1613-13656 NGC3109-8114 NGC6822-27107
M33-72475 M31-148068 M31-762645 IC1613-14055 NGC6822-3644 NGC6822-27519
M33-72577 M31-181722 M31-798387 IC1613-15060 NGC6822-5818 NGC6822-28156
M33-72755 M31-192486 M31-825408 IC1613-17477 NGC6822-6615 NGC6822-28699
M33-75218 M31-234499 M31-840950 IC1613-20305 NGC6822-9575 NGC6822-29242
M33-75973 M31-238753 SextansB-5786 IC1613-20366 NGC6822-9663 NGC6822-30436
M33-78021 M31-249436 SextansB-6281 IC1613-21316 NGC6822-9932 IC10-5584
M33-78763 M31-251868 SextansB-6394 IC1613-24234 NGC6822-11910 IC10-6521
NGC4826-59 M31-269242 SextansB-12820 IC1613-24827 NGC6822-15396 IC10-9582
NGC4826-160 M31-279262 NGC4736-1029 IC1613-25189 NGC6822-16091 IC10-43559
NGC4826-4047 M31-284554 NGC4736-5124 IC1613-26535 NGC6822-18630
NGC2403-4500 M31-295938 IC1613-251 IC1613-26813 NGC6822-19248
NGC2403-5990 M31-297792 IC1613-1015 IC1613-27793 NGC6822-19272
Since our catalog is based on IR photometry, we are particularly sensitive to dusty, evolved stars. This makes us especially effective at detecting and characterizing the population of dusty YSGs, which are likely the evolutionary descendants of luminous RSGs. The ‘RSG problem’, i.e. the observed scarcity of luminous RSGs exploding as Type-II supernovae (e.g. Smartt et al. 2009; Smartt 2015; Davies & Beasor 2020) suggests that the most luminous RSGs may evolve to warmer temperatures, either by undergoing a blue loop phase (e.g. Yoon & Cantiello 2010; Koumpia et al. 2020; Yang et al. 2023; Zapartas et al. 2024) or due to binary interaction because RSG winds are not strong enough to strip RSGs (e.g. Beasor et al. 2020; Decin et al. 2024; Antoniadis et al. 2024). Due to their high mass-loss rates during the RSG phase, these YSGs are often surrounded by significant amounts of material where dust actively forms (Gordon & Humphreys, 2019; Antoniadis et al., 2024; Decin et al., 2024).
To investigate potentially interesting candidates in our sample we considered only the YSGs fulfilling the quality criteria (as set for final probability and band completeness in Section 2.4) in all galaxies. In Fig. 6 we show the distribution of their IR color $[3.6]-[4.5]$. We fit this distribution with a Gaussian function to find a mean value of --0.07 mag (consistent with 0, which is the expected mean value) and a standard deviation of 0.24 mag. This value is also consistent with the value (0.25 mag) that has been used previously (e.g. Kourniotis et al. 2017) based on the LMC and SMC results (see Bonanos et al. 2009, 2010). Hence, we find 159 sources above the $1\sigma$ threshold (indicated in Fig. 6), which we identify as dusty YSGs (corresponding to a fraction of $\sim 0.07$ with respect to their total number of 2071 sources). It is interesting to note that there are 8 sources above the $2\sigma$ value (0.48 mag), namely M33-59527, M33-61917, M33-75218, M31-439225, M31-439242, M31-523223, WLM-19863, and NGC6822-30436), and another two (IC1613-14055 and WLM-234) above the $3\sigma$ threshold (0.72 mag). In Table 6 we present the IDs of all sources above the $1\sigma$ threshold. These are potentially Yellow Hypergiants which have suffered episodic mass loss and can provide a link to the ”RSG problem”. These sources need to be spectroscopically studied to verify their status.
5 Summary and Conclusions
In this study, we presented a comprehensive catalog of massive stars across 26 galaxies within 5 Mpc, leveraging a machine-learning classifier trained on optical and infrared photometry. Our classifier successfully classified 1,147,650 sources, of which 276,657 were deemed robust classifications based on probability and completeness criteria. Among these, we identified 120,479 RSGs, 2082 YSGs, 616 BSGs, 72 B[e] Supergiants, 150,151 WR stars, and 60 LBVs. A key result of our study is the effectiveness of the classifier across a broad metallicity range (0.07–1.36 Z), demonstrating its applicability even at low metallicities ($\sim 0.1$ Z), despite not being explicitly trained for such environments. The classifier remains robust at distances $\leq 1.5$ Mpc, with only a slight decline beyond 3 Mpc due to the spatial resolution limits of Spitzer .
We investigated the effect of metallicity on different stellar populations, finding expected trends, such as a decrease in WR stars at lower metallicities and a relative increase in BSGs and YSGs. However, a number of selection biases, among which Spitzer ’s sensitivity to dusty evolved stars, must be considered when interpreting these trends. We also identified 21 luminous RSGs ($\textrm{log}(L/L_{\odot})\,$$\geq 5.5$), including 6 extreme RSGs in M31 ($\textrm{log}(L/L_{\odot})\,$$\geq 6$), challenging the Humphreys-Davidson limit. Further investigation of these sources is necessary to confirm their nature and more accurately determine their luminosity. Additionally, 159 dusty YSGs were detected. These are optimal candidates of Yellow Hypergiants, key sources to understand the ”RSG problem”, and follow-up observations are vital.
Our catalog provides the largest sample of machine-learning-classified massive stars in nearby galaxies, making it a crucial reference for future studies. It enables the identification of prime targets for spectroscopic follow-up, particularly luminous RSGs and YHGs, to further investigate their evolutionary pathways and the role of episodic mass loss in massive star evolution. Furthermore, we compiled a catalog of 5,273 spectroscopically confirmed sources for all the 26 galaxies as derived from the literature. This offers a unique dataset of reference regarding spectral types for all massive stars and candidates known so far (including additional $\sim 330$ other sources). The accuracy of our classifier can be enhanced by incorporating additional spectral data and improving the handling of observational biases. Expanding this method with JWST photometry will allow the study of massive stars in more distant galaxies, providing deeper insights into their role in galactic evolution.
Acknowledgements. GM, AZB, KA, GMS, EC, SdW acknowledge funding support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (“ASSESS”, Grant agreement No. 772086). EZ acknowledges support from the Hellenic Foundation for Research and Innovation (H.F.R.I.) under the “3rd Call for H.F.R.I. Research Projects to support Post-Doctoral Researchers” (Project No: 7933).
Facilities: This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. The Pan-STARRS1 Surveys (PS1) and the PS1 public science archive have been made possible through contributions by the Institute for Astronomy, the University of Hawaii, the Pan-STARRS Project Office, the Max-Planck Society and its participating institutes, the Max Planck Institute for Astronomy, Heidelberg and the Max Planck Institute for Extraterrestrial Physics, Garching, The Johns Hopkins University, Durham University, the University of Edinburgh, the Queen’s University Belfast, the Harvard-Smithsonian Center for Astrophysics, the Las Cumbres Observatory Global Telescope Network Incorporated, the National Central University of Taiwan, the Space Telescope Science Institute, the National Aeronautics and Space Administration under Grant No. NNX08AR22G issued through the Planetary Science Division of the NASA Science Mission Directorate, the National Science Foundation Grant No. AST-1238877, the University of Maryland, Eotvos Lorand University (ELTE), the Los Alamos National Laboratory, and the Gordon and Betty Moore Foundation. The UHS is a partnership between the UK STFC, The University of Hawaii, The University of Arizona, Lockheed Martin and NASA.
Software: This research made use of NumPy (Harris et al., 2020), pandas (Wes McKinney, 2010; Mckinney, 2011), matplotlib (Hunter, 2007), imbalanced-learn (Lemaître et al., 2017), scikit-learn (Pedregosa et al., 2011), Astropy, a community-developed core Python package for Astronomy (Astropy Collaboration et al., 2013, 2018), Jupyter Notebooks (Kluyver et al., 2016), TOPCAT, an interactive graphical viewer and editor for tabular data (Taylor, 2005).
The NASA/IPAC Extragalactic Database (NED) is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. This research has made use of NASA’s Astrophysics Data System, and SIMBAD database, operated at CDS, Strasbourg, France.
Appendix A Classified sources from the literature
The following table contains references and the number of sources (per reference) obtained from the literature. We have taken care to remove any duplicates in the catalog, as well as to provide the most recent and precise classifications. We also excluded known foreground sources but we kept any other source that was probably member of the corresponding galaxy.
Table 7: List of references with their corresponding number of sources that contribute to our collected literature catalog.
Galaxy (Total) Reference Sources
WLM (87) Venn et al. (2003) 2
Bresolin et al. (2006) 38
Levesque & Massey (2012) 11
Britavskiy et al. (2015) 13
Britavskiy et al. (2019) 1
Maravelias et al. (2023) 1
Bonanos et al. (2024) 21
NGC 55 (279) Castro et al. (2008) 168
Patrick et al. (2017) 11
Maravelias et al. (2023) 1
Bonanos et al. (2024) 99
IC 10 (137) Massey et al. (2007b) 30
Massey et al. (2007a) 3
Tehrani et al. (2017) 29
Britavskiy et al. (2019) 6
de Wit et al. 2025 (subm.) 69
M31 (1170)* Massey et al. (2009) 2
Drout et al. (2009) 18
Neugent et al. (2012) 3
Massey et al. (2016) 951
Gordon et al. (2016) 81
Shara et al. (2016) 1
Humphreys et al. (2017) 6
Massey et al. (2019) 17
Neugent et al. (2019) 37
Kraus (2019) 11
Sholukhova et al. (2020) 2
Sarkisyan et al. (2022) 1
Neugent & Massey (2023) 40
NGC 247 (64) Solovyeva et al. (2020) 2
Maravelias et al. (2023) 1
Bonanos et al. (2024) 61
NGC 253 (81) Comerón et al. (2003a) 11
Heida et al. (2015) 1
Maravelias et al. (2023) 1
Bonanos et al. (2024) 78
NGC 300 (793) Bresolin et al. (2002) 662
Schild et al. (2003) 46
Crowther et al. (2007) 103
Gazak et al. (2015) 27
Roth et al. (2018) 504
González-Torà et al. (2022) 164
Maravelias et al. (2023) 2
Bonanos et al. (2024) 122
141414 * Using the source list from Paper I. Updated classifications for three candidate to confirmed LBVs (Sarkisyan et al. 2022; Sholukhova et al. 2020), and 11 WR stars (Neugent & Massey 2023). Removed two sources found as duplicates (original IDs: M31-957 and M31-982). Added WRs from Neugent & Massey (2023). 1 Spectral types are from Comerón et al. (2003a), while corrected coordinates from Comerón et al. (2003b). 2 Updated entries with the most recent $V$ magnitudes for 40 BSGs from Bresolin et al. (2005). 3 Updated the list of sources presented in Schild et al. (2003), by adding 10 more sources and removing 2 (as non-WR). 4 Updated spectral types for sources initially in Roth et al. (2018) (these entries were removed from the corresponding list).
Table 8: Continuation of the table.
IC 1613 (86) Kurtev et al. (2001) 1
Bresolin et al. (2007) 53
Herrero et al. (2012) 1
Garcia & Herrero (2013) 12
Britavskiy et al. (2019) 3
Chun et al. (2022) 105
de Wit et al. 2025 (subm.) 6
M33 (1547)** Massey et al. (1996) 3
Massey (1998) 52
Massey & Johnson (1998) 48 6
Urbaneja et al. (2002) 4
Bruhweiler et al. (2003) 2
Urbaneja et al. (2005) 11 7
Massey et al. (2007a) 12
Neugent & Massey (2011) 2
Drout et al. (2012) 11
Humphreys et al. (2014) 4
Massey et al. (2016) 1190
Gordon et al. (2016) 12 6
Martin & Humphreys (2017) 2
Humphreys et al. (2017) 24
Kourniotis et al. (2018) 4
Massey et al. (2019) 10 6,8
Neugent et al. (2019) 46
Kraus (2019) 7
Maryeva et al. (2020) 2 9
Smith et al. (2020) 1 9
Neugent (2021) 82 10
Liu et al. (2022) 18
Phoenix Dwarf (8) Menzies et al. (2008) 2
Saviane et al. (2009) 1
Britavskiy et al. (2015) 5
NGC 1313 (97) Hadfield & Crowther (2007) 80
Bonanos et al. (2024) 17
NGC 2366 (0) n/a 0
NGC 2403 (68) Humphreys et al. (2019) 27
Bresolin et al. (2022) 40
Kaldybekova et al. (2023) 1
M81 (82) Kudritzki et al. (2012) 26
Khan et al. (2013) 7
Humphreys et al. (2019) 28
Gómez-González et al. (2020) 21
Sextans B (2) Britavskiy et al. (2019) 2
NGC 3109 (127) Evans et al. (2007) 90
Flores-Durán et al. (2017) 1711
Davidge (2018) 3
Maravelias et al. (2023) 1
Bonanos et al. (2024) 16
NGC 3077 (0) n/a 0
151515 5 Two sources (stars 4 and 12) were removed as they provided broader classifications rather than the types in Britavskiy et al. (2019) originally. ** Using IDs from Paper I.
6 Some $V$ magnitudes retrieved from Massey et al. (2007b).
7 IDs OB 10-3 and UIT 136 refer to the same object. Kept the latest classification from that paper.
8 Some $V$ magnitudes retrieved from Massey et al. (2016).
9 Confirmed LBV candidate from Massey et al. (2016).
10 Updated an uncertain YSG from Massey et al. (2016) to a RSG. 11 $V$-band photometry from Peña et al. (2007). Hii regions 41 – 49 are knots or clumps in extended Hii regions, so they were excluded from the comparison.
Table 9: Continuation of the table.
Sextans A (143) Kaufer et al. (2004) 5
Britavskiy et al. (2014) 2
Britavskiy et al. (2015) 6
Camacho et al. (2016) 9
Garcia et al. (2019) 2
Lorenzo et al. (2022) 106
Bonanos et al. (2024) 13
NGC 4214 (0) n/a 0
NGC 4736 (3) Solovyeva et al. (2019) 2
Solovyeva et al. (2021) 112
NGC 4826 (0) n/a 0
M83 (241) Hadfield et al. (2005) 10513
Bresolin et al. (2016) 14
Della Bruna et al. (2022) 6613,14
Bonanos et al. (2024) 56
NGC 5253 (0) n/a 0
NGC 6822 (83) Massey (1998) 9
Bianchi et al. (2001) 1
Massey et al. (2007a) 2
Levesque & Massey (2012) 12
Patrick et al. (2015) 10
de Wit et al. 2025 (subm.) 49
Pegasus DIG (9) Massey et al. (2007a) 1
Britavskiy et al. (2015) 6
Britavskiy et al. (2019) 2
NGC7793 (166) Bibby & Crowther (2010) 7815
Khan et al. (2013) 3
Della Bruna et al. (2020) 5
Wofford et al. (2020) 36
Della Bruna et al. (2021) 5
Maravelias et al. (2023) 1
Bonanos et al. (2024) 38
161616 12 Magnitude derived from Solovyeva et al. (2019). 13 Assigned WR to all sources but kept the original classification within ”[]”. A magnitude value for the HeII $\lambda$4685 filter is provided wherever available else a flag value of -99 is assigned. 14 Removed WR64 as duplicate to WR44, but WR35 and WR37 were kept as different (separation at 0.999″). 15 Removed all ”not WR” objects (20), but kept ”WN?” and ”WC?” sources as WR candidates.
Appendix B Comparing literature to predicted classes
In this Appendix we provide a detailed description of the comparison results between the literature and the classifier predictions per galaxy (except for NGC 2366, NGC 3077, NGC 4214, NGC 4826, and NGC 5253, for which no sources were found). For each case we provide the number of sources found in the literature and the statistics in two distinct groups, one regarding the best (more reliable) results for sources that fulfilled the quality criteria as defined in Section 2.4, and another one for the whole set.
Regarding the comparison between the predicted class and the literature spectral type, we followed this matching algorithm: i. a predicted ”RSG” was assumed to be compatible with RSG or candidates, sources with K or M types (even broader cases such as ’early/late K/M’), carbon stars (’C-star’,’carbon’), broader labels such as ’red’ and ’cool star’, ii. ”WR” was compatible with any WR spectral type (WC, WN, etc) including candidates, iii. ”YSG” was compatible with any YSG, F/G (including ’early/late F/G’ labels), or ’warm supergiant’, iv. ”BSG” was compatible with any O/B/A [^17] The inclusion of A-type stars in this class follows the training of the method, see Paper I. (and broader classes such as ’early/mid/late B’), OB or ’Hot Supergiant’, v. ”BeBR” was compatible with any source with a B[e] spectral type or candidate, vi. ”LBV” was compatible with known and candidate LBVs, vii. ”GAL” was compatible with any label indicative of a galaxy (e.g. ’galaxy’, ’quasar’, ’AGN’, ’QSO’), but we also considered them matching with carbon stars (pointing to elliptical galaxies where older stellar populations dominate), viii. broad spectral classifications from the literature such as ’blue’ or ’hot’, emission type stars (including labels as ’Halpha star’ and ’OBem’), Hii regions and planetary nebulae (’PN’) were considered compatible with any (predicted) class of ”WR”, ”BeBR”, ”LBV”, or ”BSG”, ix. any other very coarse or doubtful case (such as ’composite’, ’cluster’, ’star’, ’SNR’, ’neb’, ’symbiotic’, ’foreground’) were considered uncertain.
Regarding the matches between our catalog and the literature source we considered 1″search radius. These matches are referred to as ’good matches’. We point here that the result of the process depends also on the accuracy of the coordinates taken from the literature.
In Table 10 we provide the first few rows of the catalog that contains all sources with spectral classification from the literature. It includes all massive stars and candidates known, as well as another $\sim 330$ sources of point sources (such as carbon stars, background galaxies, Hii regions, planetary nebulae, clusters). We provide our own ID (with a preceding ’lit-’ to differentiate from our main catalog), an IDlit that corresponds to the ID provided in the corresponding work (if present). We provide the coordinates and the spectral type. For reference, a magnitude value in a specific filter is provided, as derived from the original paper and in some cases corresponds to an indicative value (e.g. in variable sources such as in LBVs). When not available, a flagged value of -99 is given. Instead of the typical citation style we give the full bibliography code as provided by NASA’s Astrophysics Data System.
Table 10: Compiled catalog of sources with spectral classifications from the literature, including all massive stars and candidates, as well as additional point sources.
ID IDlit R.A. Dec. Spectral Type Magnitude Filter Bibcode
(deg) (deg) (mag)
lit-WLM-1 [SC85b] 15 0.498125 --15.490805 A5 Ib 18.1 V 2003AJ….126.1326V
lit-WLM-2 [SC85b] 31 0.502625 --15.475055 A5 Ib 18.4 V 2003AJ….126.1326V
lit-WLM-3 A12 0.471833 --15.477694 B9 Ia 17.98 V 2006ApJ…648.1007B
lit-WLM-4 A11 0.499875 --15.472000 O9.7 Ia 18.4 V 2006ApJ…648.1007B
lit-WLM-5 A14 0.498250 --15.490638 A2 II 18.43 V 2006ApJ…648.1007B
lit-WLM-6 A9 0.488375 --15.455027 B1.5 Ia 18.44 V 2006ApJ…648.1007B
lit-WLM-7 A16 0.491250 --15.503722 A2 Ia 18.44 V 2006ApJ…648.1007B
lit-WLM-8 A19 $0$-15.521111 G2 I 18.62 V 2006ApJ…648.1007B
lit-WLM-9 B12 0.502667 --15.474972 A2 II 18.77 V 2006ApJ…648.1007B
lit-WLM-10 B13 0.473546 --15.474972 B1 Ia 18.92 V 2006ApJ…648.1007B
 [^18] This table is available in its entirety in the VizieR/CDS catalog tool. A portion is shown here for guidance regarding its form and content.
B.1 WLM
A total of 87 sources were collected from 7 works in the literature, for which we obtained good matches (i.e. within 1″ ) for 62 sources ($\sim 71\%$). We first checked the statistics for sources with the most reliable predictions, amounting to 30 sources (almost half of the sample). We obtained an exact classification for 86.7% (26 sources) and an erroneous classification for 13.3% (4 sources). Considering the whole sample (62 sources) we got 71.0% (44 sources) as exact matches, 4.8% (3) as matches within errors, 21.0% (13) as erroneous matches, and 3.2% (2) as uncertain. By combining the correct matches we have a 75.8% success rate.
B.2 NGC 55
For NGC 55 we collected 279 sources from 4 different works. We removed 180 sources with no match at 1″  (equivalent to 64.5% of the sample). From the remaining 99 sources we did not have any reliable prediction according to the selection criteria we defined. Accounting for the whole sample we managed to obtain 14.1% (14 sources) as exact matches, 10.1% (10) as matches within errors, 73.7% (73) as erroneous matches, and 2.0% (2) as uncertain. The total success rate is 24.2%. We note here that almost all data in NGC 55 suffer from significant missing values (band completeness equal to 0.2), which means that the classifier cannot reach its top performance.
B.3 M31
As the closest and largest galaxy the number of spectroscopically confirmed objects we managed to collect summed to 1170 sources (combining 13 different works). A bit more than half of them (57.4%, 671 sources) do not have a good match. The majority of the remaining sources (382) passed the quality criteria with a 94.0% (359) being correctly predicted by the classifier, and only a 6.0% (23) classified erroneously. When considering all 499 sources the fraction of exact matches was 82.8% (413), with another 5.6% (28) predicted correctly within the error, and 11.6% (58) classified incorrectly. The total correct fraction is 88.4%. This high fractions (especially for the good candidates) are not unexpected since the classifier has been trained on (most) of these sources.
B.4 IC 10
The number of spectroscopically confirmed classified sources for IC 10 was 137 derived from five different papers. Of these sources we did not find a good match for 79 sources (57.7%). We had 11 sources passing the quality criteria resulting in 63.6% (7 sources) predicted correctly and 36.4% (4 sources) misclassified. By considering all (58) sources we got 39.7% (23 sources) as exact matches, 6.9% (4) as matches within errors, 50.0% (29) as erroneous matches, and 3.4% (2) as unclassified. By combining the correct matches we have a total success rate of 46.6%.
B.5 NGC 247
There are 64 sources in NGC 247 with secured spectra derived from three works. Out of these, 17 sources ($\sim 26.6$%) do not have a good match within 1″. There are 11 sources that fulfilled the selection criteria and we successfully predicted their class for 9 of these (81.8%), while the other two sources were erroneously classified (18.2%). The total success rate is 81.8%. By considering all (47) sources we got 44.7% (21 sources) as exact matches, 14.9% (7) as matches within errors, 27.7% (13) as erroneous matches, and additionally 12.8% (6) uncertain objects (classified originally as ”star”, ”emission objects” without any further indication of any potential stellar signal, and composite). By combining the correct matches we have a total success rate of 59.6%.
B.6 NGC 253
We were able to gather 81 classified sources from four different works. 18 sources (22.2%) were removed because of no matching with any source from our catalog. Of the remaining 63 sources only 6 satisfied the quality criteria. From these we correctly predicted 16.7% (1 source), while 50.0% (3) were erroneously classified. There were additionally 33.3% (2 sources) that were uncertain. By considering the whole sample, we got 19.0% (12 sources) as exact matches, 15.9% (10) as matches within errors, 30.2% (19) as erroneous matches, and 34.9% (22) as uncertain. The total correct matches equals 34.9%. The rather low success rate is both due to the larger number of incorrect classifications and the fact that there are many more sources in this galaxy for which their classifications are not well constrained or they refer to objects for which the classifier has not been trained (e.g. ’cluster’ or ’nebula’).
B.7 NGC 300
A total of 793 sources were collected from the literature. However, we have to remove the vast majority of them (553 sources, i.e. 69.7%) because they did not have any good match within 1″ of our catalog. Due to the lack of Pan-STARRS coverage there are no sources that can satisfy the band completeness criterion (¿0.6). Considering all (240) sources we got 10.4% (25) as exact matches, and 5.4% (13) as matches within errors, summing to 15.8%. A fraction of 80.8% (194) was predicted erroneously, while 3.3% (8) were uncertain cases. Due to the many missing features in this galaxy, the fraction of mis-prediction is high.
B.8 NGC 1313
About 100 classified sources were collected from two separate works in the literature. The majority of them (79.4%, 77 sources) lack a good match with our catalog, as well as properties above the selection criteria we set. Therefore, by considering all sources (20) we got exact matches for 35.0% (7 sources), matches within the error for 40.0% (8), erroneous predictions for 15.0% (3), and 10.0% (2) which we could not classify (uncertain). Overall, the total correct fraction is 75.0%.
B.9 IC 1613
We collected 86 sources from 7 different works. About 64% of these (55 sources) do not have a good match with any of our catalog objects. Of the remaining 32 objects, 23 are good candidates (according to the selection criteria) and we managed to get 78.3% (18 sources) as exact matches, and 21.7% (5) erroneously classified. By considering all sources we got 64.5% (20 sources) as exact matches, 3.2% (1) as a good match within the errors, and 32.3% (10) as misclassifications. The total fraction of correctly predicted sources is 67.7%.
B.10 M33
For this galaxy we have collected 1547 sources from 22 different works. About 61.3% (948) did not have any good match with our catalog. Of the remaining 599 sources, the majority (435) is of sufficient quality, and in particular 92.9% (404) were exact matches, and 7.1% (31) were erroneous. Considering all sources (599) we got 81.8% (490 sources) as exact matches, 6.2% (37) as a good match within the errors, 11.9% (71) as misclassifications, and 0.2% (1) as uncertain. The total fraction of correctly predicted sources is 88.0%.
B.11 Phoenix Dwarf
Only eight objects with spectroscopic classifications were obtained from three different works in the literature. Unfortunately, the majority of them (75.0%, i.e. 6 sources) do not have any match within the 1″  search radius. Moreover, these two remaining sources do not pass the quality criteria and they are not predicted correctly. Therefore, we got a total correct fraction of 0%, which is due to the small number statistics and the fact that the band completeness (in all of these cases) is very low (at 0.2) which means that most of the features are missing.
B.12 NGC 1313
We collected 97 sources from two different works. The majority of them (79.4%, 77 sources) do not have any good match with out catalog. Of the remaining 20 sources there are no good candidates (fulfilling the selection criteria). We had 35.0% (7 sources) as exact matches, 40.0% (8) as matches within the errors, 15.0% as misclassifications, and 10.0% (2) as uncertain. Therefore, the total correct fraction of sources was 75.0%.
B.13 NGC 2403
We identified 68 sources originating from two works (Humphreys et al. 2019; Bresolin et al. 2022), and an additional (tentative) case (Kaldybekova et al. 2023). Unfortunately, 53 ($\sim 78\%$) of these sources (including the tentative case) do not have any good match with our catalog. From the remaining 15 sources 7 were of sufficient good quality and we got 28.6% (2) as exact matches, and 71.4% (5) as erroneous. When we considered all sources we got 20.0% (3) as exact matches, 26.7% (4) as matches within the errors, and 53.3% (8) as misclassifications. The total correct fraction of sources in this case was 46.7%.
B.14 M81
A total of 82 sources were collected from the four works in the literature (including a number of AGN/galaxies and Hii regions). A good match was not found for the 67.1% (55) sources. From the remaining 27 sources, 12 passed the quality criteria, and we got 66.7% (8) as exact matches, and 33.3% (4) as erroneous. When considering all 27 sources we got 55.6% (15) as exact matches, 7.4% (2) as matches within the errors [^19] We note here that the catalog of Gómez-González et al. (2020) includes two separate sources, namely WR-3 and WR-14, which are reported as different sources with a separation of 0.87”. We opted not to remove any of them from our comparison list, although our search radius is 1″  . Therefore, these two sources are matched with the same object from our catalog (with ID M81-471) and it is a duplicate result. The impact is not significant though since this means a 3.6% decrease in the correct within the error and the total correct fraction., and 37.0% (10) as misclassifications. The total correct fraction in this case was 63.0%.
B.15 Sextans B
Spectroscopic classifications hardly exist for this galaxy. There are a few objects studied by Britavskiy et al. (2019), but the majority are foreground sources (and therefore excluded from further consideration). We were left with only two sources, of which one had a match within 1″ . This source fulfilled the section criteria and it was actually predicted correctly (leading to an obviously biased fraction of 100%).
B.16 NGC 3109
A total of 127 sources were gathered from five different works in the literature (including a number of Hii regions and planetary nebulae). A good match was not found for the majority of the sources (74.9%, 94 sources). From the remaining 33 sources, 15 passed the quality criteria, and we got 66.7% (10) as exact matches, 26.7% (4) as erroneous, and 6.7% (1) as uncertain. When considering all available sources we got 60.6% (20) as exact matches, 6.1% (2) as matches within the errors, 30.3% (10) as erroneous, and 3.0% (1) uncertain. The total correct fraction in this case was 66.7%.
B.17 Sextans A
We collected 143 classified sources from 7 different works. A significant fraction of 63.6% of them (91 sources) were found more than 1″  away from our catalog sources. There were 14 sources with quality data of which we got 64.3% (9) as exact matches, and 35.7% (5) as erroneous. By considering all 52 sources, we got 28.8% (15) as exact matches, 30.8% (16) matches within the error, and 40.4% (21) as erroneous. The total success rate is 59.6%.
B.18 NGC 4736
We only found three sources from the literature. Interestingly all of them are candidate LBVs (Solovyeva et al. 2019) of which one has been confirmed based on photometric and spectroscopic variability (Solovyeva et al. 2021). Unfortunately none of them has a good match with any of our catalog sources.
B.19 M83
We managed to obtain 241 sources from four different works. Unfortunately, the majority of them (185, 76.8%) are not matched with our catalog sources. Out of the remaining 56 objects only 8 fulfilled the selection criteria. We got 62.5% (5 sources) as exact matches, 25.0% (2) as erroneous and 12.5% (1) as uncertain. Considering all (56) objects, we got 57.1% (32) as exact matches, 14.3% (8) as correctly predicted within the error, 16.1% (9) as erroneous predictions, and 12.5% (7) as uncertain (due to the classification as clusters). The total correct fraction of our prediction was 71.4%.
B.20 NGC 6822
We obtained 83 sources from the literature (6 different works). About 31% (26 sources) did not have a good match. Of the remaining 57 sources, 31 passed the quality criteria. We actually predicted all of them (96.8%, 30 source) correctly, without any misclassification, but only 1 source (3.2%) remained as unclassified. This can be justified because the spectral types of these sources are RSG and we have the best success rate for this particular class. Considering all sources, we got 57.9% (33 sources) correctly classified, 22.8% (13) correct within the classification error, 15.8% (9) misclassified, and another 3.5% (2) as uncertain. The total correct fraction in this case was 80.7%.
B.21 Pegasus DIG
We found 9 sources from three different works. Two thirds of this sample (66.7%, 6 sources) did not have any match within 1″ . From the remaining 3 objects none fulfilled the selection criteria. Of these we got one predicted accurately, one correct within the classification error, and an erroneous one. Therefore, the success rate was at 66.7%, but it is based on a very small number of sources.
B.22 NGC 7793
We collected 166 classified sources from 7 different works. A significant fraction (65.7%, 109 sources) did not have any good match with our catalog. From the remaining 57 sources only two had quality data, and both of them were erroneously predicted (leading to a 0% success rate). Considering all sources we get 10.5% (6) as exact matches, 15.8% (9) as matches within the error, 66.7% (38) predicted erroneously, and 7.0% (4) as uncertain. The total success rate in this case was 26.3%.
Appendix C A Baeysian approach to determine uncertainties in fractions
In this Appendix, we describe the Bayesian approach used to determine the uncertainties in the fractions of the correct predictions. The likelihood is the probability of classifying correctly $k$ sources out of the $n$ of those in a certain class, given the success probability $p$ of the classifier. Consequently, it is a binomial distribution:
$P(k|n,p)=\binom{n}{k}p^{k}(1-p)^{n-k}.$ (1)
The posterior probability of the success probability is given by Bayes’ theorem:
$P(p|k,n)=\frac{P(k|n,p)P(p)}{P(k|n)},$ (2)
where $P(p)$ is the prior probability of the success rate, and $P(k|n)$ is the marginal likelihood which can be treated as a normalization constant since it does not depend on $p$.
In this work, we report as point estimate the mode of the posterior, $m=\text{arg}\max\limits_{p}P(p|k,n)$. Allowing for custom priors, we perform the calculation numerically with a resolution of 0.001 in $p$, which is well below the resulting uncertainties by the range of values for $n$ and $k$ in our sample.
The reported uncertainty corresponds to the 68% highest posterior density interval (HPDI), i.e., the smallest interval that contains 68% of the posterior probability:
HPDI $=\left[l,u\right],$ (3)
[…]
where the lower, $l$, and upper, $u$, bounds are such that they contain the mode, and have equal probability density except for cases where the interval includes the extrema of $p$, 0 or 1, in which it can only expand towards one tail. In our implementation, the HPDI is calculated by starting from the mode, and expanding the interval at lower or higher values of $p$, depending on the direction at which the posterior density is higher and the interval can expand if it reached one of the extrema.
For the prior we explore the uniform distribution, $P(p)=1$, as well as the Beta distribution which is the conjugate distribution of the binomial, and is often employed as prior on probabilities. For the latter case, since we aim at using a prior with location and spread representing previous results, we constructed an algorithm that finds the parameters $\alpha$ and $\beta$ of the Beta distribution that result in the desired mode and variance. We do this by numerically solving for the parameters in the system of equations:
mode $=\frac{\alpha-1}{\alpha+\beta-2},$ (5)
variance $=\frac{\alpha\beta}{(\alpha+\beta)^{2}(\alpha+\beta+1)},$ (6)
where the variance is less than $\frac{1}{12}$ to ensure unimodality.
The code for the Bayesian analysis, and the search for the unimodal Beta distribution, as well as the documentiation, are provided in the following open source repository: https://github.com/kkovlakas/gaussfree.